if tom gives maria 30cents, they will have equal amount of money. but if maria gives tom 50 cents, he will have twice as much moneh as she does. how much money does each of them have?
Answer by ankor@dixie-net.com(22740)

if tom gives maria 30cents, they will have equal amount of money.
but if maria gives tom 50 cents, he will have twice as much money as she does. how much money does each of them have?
:
Let t = amt of money Tom has originally (in cents)
Let m = amt of that Maria has
:
"if tom gives maria 30cents, they will have equal amount of money."
t - 30 = m + 30
t = m + 30 + 30
t = m + 60
:
"but if maria gives tom 50 cents, he will have twice as much moneh as she does."
t + 50 = 2(m-50)
t + 50 = 2m - 100
t = 2m - 100 - 50
t = 2m - 150
replace t with (m+60)
m + 60 = 2m - 150
60 + 150 = 2m - m
210 = m
therefore, in dollars:
Maria has $2.10
then
Tom has 2.10 + .60 = $2.70
:
:
Check solutions in the statement:
maria gives tom 50 cents, he will have twice as much money as she does. "
2.70 + .50 = 2(2.10-.50)
3.20 = 2(1.60)